OpenStudy (vlery):

I just want to be sure if this is correct ( this is trig math if you're not in it or taken it don't answer it)

OpenStudy (vlery):

would this be the Sum-Difference Formula?

jimthompson5910 (jim_thompson5910):

hints: sin^2 + cos^2 = 1 -(sin^2 + cos^2) = -1 and csc = 1/sin

OpenStudy (vlery):

my teacher didn't go over the identities with us except for the Pythagorean Identities

jimthompson5910 (jim_thompson5910):

so sin^2 + cos^2 = 1 looks familiar right?

OpenStudy (vlery):

ohhhh haha so my equation is a Pythagorean Identity?

jimthompson5910 (jim_thompson5910):

also, does \[\Large \csc(\theta) = \frac{1}{\sin(\theta)}\] look familiar?

OpenStudy (vlery):

that's the Reciprocal identity

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

so use that idea to get \[\Large \sin(\theta)\csc(\theta) - \sin^2(\theta) = \cos^2(\theta)\] \[\Large \sin(\theta)*\frac{1}{\sin(\theta)} - \sin^2(\theta) = \cos^2(\theta)\] what's next?

OpenStudy (vlery):

\[1/\sin(\theta)\sin(\theta)\] on the right side?

jimthompson5910 (jim_thompson5910):

when it comes to proving identities, you only work on one side and leave the other side alone

jimthompson5910 (jim_thompson5910):

it's easier to start with a more complicated side and simplify it down so that's why I'm starting on the left side

OpenStudy (vlery):

would you do something about that \[\sin^2\]

jimthompson5910 (jim_thompson5910):

what happens with the sin times 1/sin

OpenStudy (vlery):

ohhhhhh this might take a bit but I think I got it

OpenStudy (vlery):

 nvm that's all I got